# CLAPSep: Leveraging Contrastive Pre-trained Model for Multi-Modal Query-Conditioned Target Sound Extraction

### v3

@@ V-D Practical Scenarios Analysis and Evaluation
When evaluating the model’s robustness in dealing with different sound events, we construct synthetic sound mixtures with 2, 3, and 4 sound sources and extract 1, 2, and 3 target sound events simultaneously. Corresponding results are shown in Table IV, where CLAPSep-P denotes CLAPSep with positive language query and CLAPSep-PN denotes CLAPSep with positive and negative language queries. Though we construct the training audio mixture using one target sound with one interference noise, the model’s capability in handling multiple sound sources is made possible thanks to the complexity of the sound samples in the AudioCaps dataset, i.e., a single audio clip often contains more than one sound event. We can simply write a prompt that “The sounds of A and B.” to extract sound events A and B simultaneously. As can be seen from the results in Table IV, our proposed method has consistent and significant performance gains relative to existing methods when dealing with multiple sound sources.
When evaluating the model’s performance in dealing with different input SDRs, we synthesize evaluation sound mixtures with SDR varying from -10 to 10 dB. As in our initial training setting, the training sound mixtures are only constructed at 0 dB SDR, varying the input SDR in the testing stage is out of distribution thus we find the performance drops significantly as indicated in row 2 on Tabel V. To mitigate this, We simply fine-tune the pre-trained model on sound mixtures constructed on SDR that uniformly distributed in the -10 to 10 dB interval, leaving all settings unchanged except for decreasing the learning rate to 1e-5. We fine-tune a total of 5 epochs and the results are shown in the third row of Table V denoted as CLAPSep. It can be seen that the fine-tuned model is more robust to a variety of input SDRs and outperforms AudioSep consistently, despite a slight performance degradation at 0 dB. It is worth noting that the AudioSep model we use was originally trained on audio mixtures constructed on uniformly distributed SDR in the range of -10 to 10 dB, and therefore we do not fine-tune the AudioSep model.
@@ V-E Ablation Study
This section investigates the impact of each designed component in CLAPSep. The ablation study is carried out under four settings as shown in Table VI: 1) w/o CLAP audio encoder, where we discard the CLAP audio encoder for extracting layer-wise audio features. 2) w/o pre-trained weights, where the audio encoder is randomly initialized and trained from scratch. 3) frozen pre-trained weights, where parameters of the audio encoder are frozen during training. 4) replace text encoder to CLIP [26], where the CLAP text encoder in the query network is replaced with CLIP text encoder. All other conditions, such as training data and the number of training steps, are held constant to ensure a fair comparison.
As shown in Table VI, removing the CLAP audio encoder (w/o CLAP audio enc.) leads to a noticeable reduction of performance, confirming the value of both the hierarchical structure of the encoder and the pre-trained weights. Similarly, replacing the pre-trained encoder in the separation network with a randomly initialized one (w/o pre-trained weights) results in a substantial performance drop, particularly in cross-dataset evaluations. This decline emphasizes the importance of the embedded prior knowledge within the pre-trained CLAP, highlighting its role in enhancing the model’s generalizability. Additionally, the exclusion of LoRA (frozen weights) slightly decreases performance, underscoring its effectiveness in fine-tuning the pre-trained encoder for specific sound extraction tasks. We also investigate how different language query networks influence the overall system performance by replacing the CLAP text encoder in the query network with a pre-trained CLIP text encoder (replace text encoder to CLIP). This replacement causes a certain degree of performance decline, particularly in cross-dataset evaluations. This demonstrates that our proposed method, through tightly coupled query and separation networks, has elicited the model’s optimal performance and generalizability. This is attributed to the similar structure and shared semantic space of the separation network and the CLAP text encoder. Notably, such a replacement for the query network causes almost no performance loss for AudioSep [21] where the separation network and query encoders are not semantically and structurally related. Fig. 7 describes the training dynamics of CLAPSep under different setups. It can be observed that the CLAP audio encoder accelerates the training process and achieves better performance due to both its advanced model architecture and the encoded prior knowledge in pre-training.
@@ V-F Automatic Negative Query Embedding Generation
TABLE VII: Evaluations w/ or w/o negative embedding generation (Algorithm 1).
@@
Table VII presents the experimental results on MUSIC21, ESC-50, and FSDKaggle2018. By integrating the negative embeddings generated by the proposed Algorithm 1, denoted P+N (Alg 1), there is a notable performance improvement over the positive-only querying (P), especially on the MUSIC21 dataset. This demonstrates the effectiveness of CLAPSep for automatic negative query embedding generation. We further design experiments to verify how the selection and aggregation of pre-cached embeddings (as described in step 4) affects the system performance. Omitting this step (referred to as skip Step 4) leads to a performance decline. This decline may be attributed to the fact that the extracted sound embedding $\hat{\bm{e}}^{N}$ retains some information about the target sound. Therefore, step 4 is crucial as it aligns the sound embedding with cached negative query embeddings, effectively purging this residual information and resulting in a purified negative query embedding.
@@ V-G Visualization Analysis
The model’s proficiency in target sound extraction and non-target sound suppression is evident in the second and third rows, where positive and negative query captions are the sole inputs. The model can effectively extract target sound sources or suppress non-target sound sources when only a positive or negative user query is provided. The fourth and fifth rows highlight environmental sound extraction queried by labels (which are transformed to language descriptions by adding the prefix “The sound of”) and instrumental music extraction queried by audio samples. A comparison of the spectrograms for sound mixtures, separated sources, and ground truth reveals that CLAPSep adeptly preserves the desired sound sources while eliminating unwanted ones. This demonstrates the effective handling of multi-modal clues queried target sound extraction tasks by CLAPSep.
@@ VI Limitations Discussion
While our proposed method advances the field of query-conditioned target sound extraction, certain limitations warrant further exploration. Firstly, our CLAPSep model is non-causal, meaning that the separation outcome at any given time step relies on both preceding and subsequent input frames. This dependency restricts its applicability in real-time streaming applications. Secondly, CLAPSep uses the phase of the sound mixture directly as the phase estimate for the target sound during the inverse short-time Fourier transform. Future enhancements could include the use of neural networks to estimate phase residuals, potentially leading to further improvements in performance.
@@ VII Contributions and Conclusion
This paper proposes CLAPSep for target sound extraction. CLAPSep extracts the target sound from sound mixtures, responding to versatile, multi-modal user queries, which can be text-only, audio-only, or a combination of both. Notably, it supports both positive and negative queries to facilitate fine-grained sound event descriptions, thus further enhancing its sound extraction capabilities. The effectiveness of CLAPSep is primarily due to two key factors: the ability to handle fine-grained, diverse user queries and the leverage of the advanced CLAP model at its core. Extensive qualitative and quantitative experiments conducted across 5 benchmarks show that CLAPSep provides significant and consistent performance improvements over prior methods. In addition, CLAPSep demonstrates decent zero-shot generalizability to previously unseen sound events, suggesting it could be a significant step forward in universal sound separation technology. To support ongoing research and encourage further development, we have made the code and model checkpoints publicly available.

### v5

@@ V-D Practical Scenarios Analysis and Evaluation
When evaluating the model’s robustness in dealing with different sound events, we construct synthetic sound mixtures with 2, 3, and 4 sound sources and extract 1, 2, and 3 target sound events simultaneously. Corresponding results are shown in Table V, where CLAPSep-P denotes CLAPSep with positive language query and CLAPSep-PN denotes CLAPSep with positive and negative language queries. Though we construct the training audio mixture using one target sound with one interference noise, the model’s capability in handling multiple sound sources is made possible thanks to the complexity of the sound samples in the AudioCaps dataset, i.e., a single audio clip often contains more than one sound event. We can simply write a prompt that “The sounds of A and B.” to extract sound events A and B simultaneously. As can be seen from the results in Table V, our proposed method has consistent and significant performance gains relative to existing methods when dealing with multiple sound sources.
When evaluating the model’s performance in dealing with different input SDRs, we synthesize evaluation sound mixtures with SDR varying from -10 to 10 dB. As in our initial training setting, the training sound mixtures are only constructed at 0 dB SDR, varying the input SDR in the testing stage is out of distribution thus we find the performance drops significantly as indicated in row 2 on Tabel VI. To mitigate this, We simply fine-tune the pre-trained model on sound mixtures constructed on SDR that uniformly distributed in the -10 to 10 dB interval, leaving all settings unchanged except for decreasing the learning rate to 1e-5. We fine-tune a total of 5 epochs and the results are shown in the third row of Table VI denoted as CLAPSep. It can be seen that the fine-tuned model is more robust to a variety of input SDRs and outperforms AudioSep consistently, despite a slight performance degradation at 0 dB. It is worth noting that the AudioSep model we use was originally trained on audio mixtures constructed on uniformly distributed SDR in the range of -10 to 10 dB, and therefore we do not fine-tune the AudioSep model.
@@ V-E Ablation Study
This section investigates the impact of each designed component in CLAPSep. The ablation study is carried out under four settings as shown in Table VII: 1) w/o CLAP audio encoder, where we discard the CLAP audio encoder for extracting layer-wise audio features. 2) w/o pre-trained weights, where the audio encoder is randomly initialized and trained from scratch. 3) frozen pre-trained weights, where parameters of the audio encoder are frozen during training. 4) replace text encoder to CLIP [26], where the CLAP text encoder in the query network is replaced with CLIP text encoder. All other conditions, such as training data and the number of training steps, are held constant to ensure a fair comparison.
As shown in Table VII, removing the CLAP audio encoder (w/o CLAP audio enc.) leads to a noticeable reduction of performance, confirming the value of both the hierarchical structure of the encoder and the pre-trained weights. Similarly, replacing the pre-trained encoder in the separation network with a randomly initialized one (w/o pre-trained weights) results in a substantial performance drop, particularly in cross-dataset evaluations. This decline emphasizes the importance of the embedded prior knowledge within the pre-trained CLAP, highlighting its role in enhancing the model’s generalizability. Additionally, the exclusion of LoRA (frozen weights) slightly decreases performance, underscoring its effectiveness in fine-tuning the pre-trained encoder for specific sound extraction tasks. We also investigate how different language query networks influence the overall system performance by replacing the CLAP text encoder in the query network with a pre-trained CLIP text encoder (replace text encoder to CLIP). This replacement causes a certain degree of performance decline, particularly in cross-dataset evaluations. This demonstrates that our proposed method, through tightly coupled query and separation networks, has elicited the model’s optimal performance and generalizability. This is attributed to the similar structure and shared semantic space of the separation network and the CLAP text encoder. Notably, such a replacement for the query network causes almost no performance loss for AudioSep [21] where the separation network and query encoders do not establish a prior semantic and structural association. Fig. 7 describes the training dynamics of CLAPSep under different setups. It can be observed that the CLAP audio encoder accelerates the training process and achieves better performance due to both its advanced model architecture and the encoded prior knowledge in pre-training.
@@ V-F Automatic Negative Query Embedding Generation
TABLE VIII: Evaluations w/ or w/o negative embedding generation (Algorithm 1).
@@
Table VIII presents the experimental results on MUSIC21, ESC-50, and FSDKaggle2018. By integrating the negative embeddings generated by the proposed Algorithm 1, denoted P+N (Alg 1), there is a notable performance improvement over the positive-only querying (P), especially on the MUSIC21 dataset. This demonstrates the effectiveness of CLAPSep for automatic negative query embedding generation. We further design experiments to verify how the selection and aggregation of pre-cached embeddings (as described in step 4) affects the system performance. Omitting this step (referred to as skip Step 4) leads to a performance decline. This decline may be attributed to the fact that the extracted sound embedding $\hat{\bm{e}}^{N}$ retains some information about the target sound. Therefore, step 4 is crucial as it aligns the sound embedding with cached negative query embeddings, effectively purging this residual information and resulting in a purified negative query embedding.
@@ V-G Visualization Analysis
The model’s proficiency in target sound extraction and non-target sound suppression is evident in the second and third rows, where positive and negative query captions are the sole inputs. The model can effectively extract target sound sources or suppress non-target sound sources when only a positive or negative user query is provided. The fourth and fifth rows highlight environmental sound extraction queried by labels (which are transformed to language descriptions by adding the prefix “The sound of”) and instrumental music extraction queried by audio samples. A comparison of the spectrograms for sound mixtures, separated sources, and ground truth reveals that CLAPSep adeptly preserves the desired sound sources while eliminating unwanted ones. This demonstrates the effective handling of multi-modal cues queried target sound extraction tasks by CLAPSep.
@@ VI Limitations Discussion
While our proposed method advances the field of query-conditioned target sound extraction, certain limitations warrant further exploration. Firstly, our CLAPSep model is non-causal, meaning that the separation outcome at any given time step relies on both preceding and subsequent input frames. This dependency restricts its applicability in real-time streaming applications. Secondly, CLAPSep uses the phase of the sound mixture directly as the phase estimate for the target sound during the inverse short-time Fourier transform. Future enhancements could include the use of neural networks to estimate phase residuals, potentially leading to further improvements in performance. Additionally, although our method leverages the prior knowledge of the pre-trained CLAP model for efficient training of the TSE model, it still relies on manually labeled text-audio pairs during training. Further exploration into language-queried target sound extraction without parallel text-audio data could unlock the potential of vast amounts of unlabeled audio for training TSE models effectively.
@@ VII Contributions and Conclusion
This paper proposes CLAPSep for target sound extraction. CLAPSep extracts the target sound from sound mixtures, responding to versatile, multi-modal user queries, which can be text-only, audio-only, or a combination of both. Notably, it supports both positive and negative queries to facilitate fine-grained sound event descriptions, thus further enhancing its sound extraction capabilities. The effectiveness of CLAPSep is primarily due to two key factors: the ability to handle fine-grained, diverse user queries and the leverage of the advanced CLAP model at its core. Extensive qualitative and quantitative experiments conducted across 5 benchmarks show that CLAPSep provides significant and consistent performance improvements over prior methods. In addition, CLAPSep demonstrates decent zero-shot generalizability to previously unseen sound events, suggesting it could be a significant step forward in universal sound separation technology. Notably, such descent performance gains are achieved with reduced data and computational demands for training relative to a series of recent TSE works [21, 9, 24]. This is evident in that the proposed method requires less than 3% of the training data and steps to achieve superior model performance and generalizability compared to the previous state-of-the-art, AudioSep. In contrast, when the same amount of training data is provided to AudioSep, its performance lags significantly behind ours. The results of the ablation experiments further indicate that the efficiency of model training is largely due to the comprehensive utilization of the prior knowledge embedded in the pre-trained CLAP model. To support ongoing research and encourage further development, we have made the code and model checkpoints publicly available.